Three capacitors with capacitances 4 F, and 12 F are connected in series. A voltage of 400V is applied to the combination.Find the charge in each capacitor.(IN C)

1
Expert's answer
2022-03-14T13:06:46-0400

When capacitors was combined in series, charge on each capacitor is equal total charge. Find total capacity in series


"C=\frac{C_1C_2}{C_1+C_2}"

For total charge


"Q=CU"

Answer 1200 C

Calculating


"Q=3\cdot400=1200""C=\frac{4\cdot12}{4+12}=3"